# Holographic Tensor Networks as Tessellations of Geometry

### v1

@@ Abstract
Holographic tensor networks serve as toy models for the Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence, capturing many of its essential features in a concrete manner. However, existing holographic tensor network models remain far from a complete theory of quantum gravity. A key obstacle is their discrete structure, which only approximates the semi-classical geometry of gravity in a qualitative sense. In [1], it was shown that a network of partial-entanglement-entropy (PEE) threads, which are bulk geodesics with a specific density distribution, generates a perfect tessellation of AdS space. Moreover, such PEE-network tessellations can be constructed for more general geometries using the Crofton formula. In this paper, we assign a quantum state to each vertex in the PEE network and develop two holographic tensor network models: the factorized PEE tensor network, which takes the form of a tensor product of EPR pairs, and the random PEE tensor network. In both models we reproduce the exact Ryu–Takayanagi formula by showing that the minimal number of cuts along a homologous surface in the network exactly computes the area of this surface.
@@
Introduction— The Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence [2, 3, 4], also known as the holographic principle, is a profound framework in theoretical physics that establishes an equivalence between a gravitational theory in a higher-dimensional spacetime and a quantum field theory on its lower-dimensional boundary. It provides a non-perturbative definition of quantum gravity in AdS space in terms of a well-defined quantum field theory without gravity. Moreover, the duality has deep connections to quantum information theory. For example, concepts like entanglement wedge reconstruction (or subregion duality) [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] reveal how quantum information in a bulk region is encoded in the boundary theory, leading to natural error-correcting properties [16, 14, 17, 18, 19, 20, 21].
@@
The tensor network approach [22, 23, 24, 25, 26, 27, 28, 29, 30, 31] provides a concrete framework to explore the structure of the AdS/CFT correspondence 111see [46, 33, 70, 24, 31, 48, 49] for an incomplete list of recent progress.. More explicitly, we start with a space filled with tensors representing quantum states on each site in the space, then we contract the “legs” (indices) to connect adjacent tensors thus get a tensor network, which plays the role of the background geometry. The uncontracted legs of tensors in the bulk and the boundary are regarded as the bulk and the boundary degrees of freedom respectively, and the tensor network provides a quantum-error-correcting code that encodes the bulk states into the boundary Hilbert space [14, 33]. Many toy models of tensor network to simulate the AdS/CFT correspondence have been proposed (see also [34, 35, 36, 37, 38] for recent interesting developments), based on different structures of the networks and different quantum states associated to the tensors. One important holographic feature for these toy models is that, the entanglement entropy is given by the minimal number of cuts along a homologous surface in the network, which matches the area of the minimal surface of the celebrated Ryu-Takayanagi (RT) formula [39, 40, 41, 42, 43, 44].
@@
Nevertheless, in most of these toy models their structures of the tensor network are usually discrete and only capture the features of the background geometry in a qualitative way. There is a big gap between number of cuts in a discrete network and the area of a surface in a smooth Riemannian manifold. Recently, a continuous network as a perfect tessellation of the background AdS geometry was constructed using the so-called partial-entanglement-entropy (PEE) threads [45, 1], which are bulk geodesics with a particular density distribution determined by the PEE structure of the boundary CFT. In such a network, the number of cuts along a surface exactly reproduces the area of this surface. In this work, we will develop two toy models of tensor network based on the PEE threads (or the PEE tensor network for short) in $d$-dimensional Poincaré AdS space. The first one is the factorized PEE tensor network constructed via tensor product of EPR pairs. It enjoys similar properties to the Harlow-Pastawski-Preskill-Yoshida (HaPPY) code [46] which is built from perfect tensors, and perfectly reproduces the RT formula for spherical regions in the vacuum CFT state. In the second model, following [33, 28, 47, 48, 49, 37] we associate a random state to each bulk site to build the random PEE tensor network, where we can exactly reproduce the RT formula for generic boundary regions.
@@
The network of PEE threads— The partial entanglement entropy (PEE) [50, 51, 52, 53, 54, 55] is a measure of entanglement $\mathcal{I}(A,B)$ between any two non-overlapping regions $A$ and $B$ featured by the key property of additivity. It can be determined by the following physical requirements [51, 56]:
@@
Here the regions are all non-overlapping, $A\cup\bar{A}$ makes a pure state and $S_{A}$ denotes the entanglement entropy of $A$. The solutions for these requirements were first studied in [41] as the mutual information that is additive in some special theories, and then the solutions are understood as the PEE in [51], which is in general not a mutual information. Later it was gradually realized that, the requirement of normalization should not be imposed for an arbitrary region. In the case of the vacuum CFT state on a plane, the solution only exists if we impose the normalization requirement to spherical regions [51]. See also [56, 50, 57, 51, 55] for other prescriptions to construct the PEE which satisfy the above requirements, and see [58, 59, 60, 61, 62, 63, 64, 65] for related discussions on PEE.
Due to the properties of additivity and permutation symmetry, any PEE $\mathcal{I}(A,B)$ can be further decomposed into a class of PEE between two different sites $\mathbf{x}$ and $\mathbf{y}$, or the two-point PEE $\mathcal{I}\left(\mathbf{x},\mathbf{y}\right)$, in the following way [51, 45]:
@@
Moreover, in the vacuum CFTd on a plane the two-point PEE can be fully determined by imposing all the above requirements for the spherical regions [45, 1], i.e.
@@
In the context of AdS/CFT, the vacuum CFTd state on a plane is dual to the Poincaré AdSd+1 space. The two-point PEE $\mathcal{I}\left(\mathbf{x},\mathbf{y}\right)$ was also proposed to have a natural holographic dual [45, 1], which is the density of the bulk geodesics connecting the boundary sites $\mathbf{x}$ and $\mathbf{y}$. The bulk geodesics whose density are determined in such a way are denoted as the PEE threads. Let us set that each PEE thread gives unit contribution to the PEE, then (1) captures the number of PEE threads connecting the two boundary regions $A$ and $B$. The collection of all the PEE threads forms a continuous, dense network permeating the entire time slice of the Poincaré AdSd+1, which we call the PEE network (see an example of $d=2$ in Fig.1). Remarkably, it was found that [45, 1] the PEE network is a tessellation of the static AdS space in the following sense:
@@
One profound implication of the above prescription is that, minimizing the area of a surface $\Sigma_{A}$ homologous to any boundary region $A$ is equivalent to minimizing the number of intersections between $\Sigma_{A}$ (see the red line in Fig.1) and the PEE network. This further implies that we can exactly reproduce the Ryu-Takayanagi (RT) formula for holographic entanglement entropy by developing proper tensor network models where the entanglement entropy of a boundary region is captured by the minimal number of intersections between a homologous surface and the PEE network [45, 1, 64]. Briefly speaking, the RT formula for the boundary region $A$ can be reformulated as follows,
@@
Figure 1: Extracted from [64]. The PEE network as a tessellation of a time slice of AdS3. Here the blue lines represent the PEE threads, the red line is a surface $\Sigma_{A}$ homologous to the boundary region AA.
Geometric setup for bulk tensors— In a model of tensor network, a vertex at site $\mathbf{z}$ with a set of legs corresponds to a quantum state represented by a local tensor $\mathcal{T}_{u_{1}u_{2}\cdots}(\mathbf{z})$, where the index $u_{i}$ characterizes the leg that is associated to a Hilbert space $\mathcal{H}_{i}$ spanned by the basis vectors $\{|u_{i}\rangle\}$. Then the quantum state in the Hilbert space $\bigotimes_{i}\mathcal{H}_{i}$ settled at $\mathbf{z}$ is defined by (assuming Einstein summation hereafter),
@@
Figure 2: A tensor state located at the bulk site $\mathbf{z}$ (hollow white dot in the center). The legs connecting $\mathbf{z}$ is actually continuously distributed, which can be described by a vector field as in [45]. Only six legs are exhibited, while other legs are collectively represented by the blue region. The contraction between tensors at $\mathbf{z}$ and at adjacent site $\mathbf{z^{\prime}}$ is realized by projecting the total state into a maximally entangled state $|\mathbf{z}\mathbf{z^{\prime}}\rangle$ along the PEE thread in between.
@@
Factorized PEE tensor network— Let us consider a pair of PEE threads which makes a complete geodesic passing through $\mathbf{z}$, and we denote it as $C_{\mathbf{z}}(\mathbf{x},\mathbf{x^{\prime}})\equiv C_{\mathbf{z}}(\mathbf{x})\cup C_{\mathbf{z}}(\mathbf{x^{\prime}})$. Correspondingly we can divide the legs of $\mathbf{z}$ into pairs and rewrite the quantum state (8) as,
@@
It has been demonstrated that [1], given a static boundary interval $A$ (or any spherical region in higher dimensions) and the corresponding RT surface $\gamma_{A}$, the PEE threads in $\mathcal{C}^{A\bar{A}}$ are the only threads that intersect with $\gamma_{A}$, and each thread only intersects once. This means the number of intersections between $\gamma_{A}$ and the PEE network is exactly $\mathcal{I}(A,\bar{A})$. Since the area of $\gamma_{A}$ can be computed by counting the number of intersections (3), we have
@@
which reproduces the RT formula for any static boundary single intervals and spherical regions. See the appendix for more details for the bulk geometric picture, and another derivation for the RT formula by constructing isometric mappings from $\gamma_{A}$ to the boundary.
Nevertheless, for disconnected and non-spherical connected boundary regions, the homologous surfaces $\Sigma_{A}$ that only intersect with the threads in $\mathcal{C}^{A\bar{A}}$ no longer exists. Furthermore, the RT surface $\gamma_{A}$ intersects not only with all the threads in $\mathcal{C}^{A\bar{A}}$, but also with a subset of the threads in $\mathcal{C}^{A}$ or $\mathcal{C}^{\bar{A}}$ [1], which means $\mathcal{I}(A,\bar{A})<\frac{Area(\gamma_{A})}{4G}$. For these reasons we fail to reproduce the RT formula for generic boundary regions in this model.
Random PEE tensor network— In the second model, we also start from the PEE network, but set the states $|\mathcal{T}(\mathbf{z})\rangle$ (8) on the bulk vertices to be unit random states following [33]. We will take the random average of such states when computing the entropy quantities. To be specific, we take an arbitrary reference state $|0_{\mathbf{z}}\rangle$ and define $|\mathcal{T}(\mathbf{z})\rangle=U_{\mathbf{z}}|0_{\mathbf{z}}\rangle$ where $U_{\mathbf{z}}$ is a unitary operator, then the random average is equivalent to an integration over $U_{\mathbf{z}}$ according to the Haar probability measure. Again, after contracting all the internal legs of adjacent bulk sites along all the PEE threads using (9), we obtain the random PEE tensor network representing a boundary state for the boundary open legs.
In [33], the authors have shown that in a general graph of network with random tensors on the vertices, if we take the large $v$ limit for the bond dimension of legs, the entanglement entropy for an arbitrary boundary region $A$ is dominant by the area term:
@@
Discussions— In this work, we assign quantum states to the vertices in the PEE network of AdS space and build two PEE tensor network models, where the entanglement entropy can be explicitly computed. In our models, counting the minimal number of cuts in the tensor network is exactly computing the area of the RT surface in the semi-classical geometry of gravity. This provides the first example of taking natural continuous limit for a discrete tensor network to perfectly match the geometric background.
In the factorized model, the RT formula is only understood for spherical regions. This is not surprising as in this model all the PEE threads are decoupled from each other, which means the model is oversimplified for AdS/CFT. The factorized model reminds us of the bit thread configurations [67, 68], and we recommend the readers to consult [45] for their similarities and differences  [^2] In [45] it was demonstrated that bit threads configuration for spherical regions $A$ can be reproduced by the superposition of the vector fields describing the PEE threads emanating from $A$..
More interesting models can be developed by including coupling among the PEE threads. In the appendix, besides regarding the tensors as tensor product of two-index tensors, we also construct a HaPPY-like [46] PEE tensor network, where we associate a perfect tensor to each site in the PEE network. The RT formula for connected regions in AdS3 is derived using a similar greedy algorithm developed in [46] for our HaPPY-like PEE tensor network.
As in [33, 46], we can also set open legs for the bulk sites to built a holographic code mapping the bulk states to the boundary states, then we can study the bulk quantum correction to the holographic entanglement entropy and identify the analogue of the quantum extremal surface [44]. Perhaps the most interesting future direction is to generalize the PEE tensor network to geometric background beyond the Poincaré AdS space, thus construct a new framework to study holographic correspondence beyond AdS/CFT. This is promising because the Crofton formula works for more generic Riemann manifolds, and we can build the network of all the geodesics in the Kinematic Space. The generalization will be non-trivial, as it is quite possible that the geodesics in the bulk will not always anchor on the boundary, which may correspond to spaces containing black holes or singularities.
Acknowledgments—The authors are supported by the NSFC Grant No. 12447108 and the Shing-Tung Yau Center of Southeast University. We thank Debarshi Basu for helpful discussions and related collaborations.

### v2

@@ Abstract
Holographic tensor networks serve as toy models for the Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence, capturing many of its essential features in a concrete manner. However, existing holographic tensor network models remain far from a complete theory of quantum gravity. A key obstacle is their discrete structure, which only approximates the semi-classical geometry of gravity in a qualitative sense. In Lin:2024dho , it was shown that a network of partial-entanglement-entropy (PEE) threads, which are bulk geodesics with a specific density distribution, generates a perfect tessellation of AdS space. Moreover, such PEE-network tessellations can be constructed for more general geometries using the Crofton formula. In this paper, we assign a quantum state to each vertex in the PEE network and develop two holographic tensor network models: the factorized PEE tensor network, which takes the form of a tensor product of EPR pairs, and the random PEE tensor network. In both models we reproduce the exact Ryu–Takayanagi formula by showing that the minimal number of cuts along a homologous surface in the network exactly computes the area of this surface.
1 Introduction
@@
The Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence Witten:1998qj ; Gubser:1998bc ; Maldacena:1997re , also known as the holographic principle, is a profound framework in theoretical physics that establishes an equivalence between a gravitational theory in a higher-dimensional spacetime and a quantum field theory on its lower-dimensional boundary. It provides a non-perturbative definition of quantum gravity in AdS space in terms of a well-defined quantum field theory without gravity. Moreover, the duality has deep connections to quantum information theory. For example, concepts like entanglement wedge reconstruction (or subregion duality) Czech:2012bh ; Hamilton:2006az ; Morrison:2014jha ; Bousso:2012sj ; Bousso:2012mh ; Hubeny:2012wa ; Wall:2012uf ; Headrick:2014cta ; Jafferis:2015del ; Dong:2016eik ; Cotler:2017erl reveal how quantum information in a bulk region is encoded in the boundary theory, leading to natural error-correcting properties Almheiri_2015 ; Dong:2016eik ; Kamal:2019skn ; Kang:2019dfi ; Akers:2021fut ; Xu:2024xbz ; Crann:2024gkv .
@@
The tensor network approach Swingle:2009bg ; Swingle:2012wq ; Qi:2013caa ; Miyaji:2015fia ; Czech:2015kbp ; Harlow:2016vwg ; Vasseur:2018gfy ; Dong:2018seb ; Hung:2019zsk ; Bao:2019fpq provides a concrete framework to explore the structure of the AdS/CFT correspondence 111see Pastawski:2015qua ; Hayden:2016cfa ; Yang:2015uoa ; Qi:2013caa ; Bao:2019fpq ; Cheng:2022ori ; Dong:2023kyr for an incomplete list of recent progress.. More explicitly, we start with a space filled with tensors representing quantum states on each site in the space, then we contract the “legs” (indices) to connect adjacent tensors thus get a tensor network, which plays the role of the background geometry. The uncontracted legs of tensors in the bulk and the boundary are regarded as the bulk and the boundary degrees of freedom respectively, and the tensor network provides a quantum-error-correcting code that encodes the bulk states into the boundary Hilbert space Dong:2016eik ; Hayden:2016cfa . Many toy models of tensor network to simulate the AdS/CFT correspondence have been proposed (see also Chandra:2023dgq ; Chen:2024unp ; Bao:2024ixc ; Kaya:2025vof ; Akella:2025owv for recent interesting developments), based on different structures of the networks and different quantum states associated to the tensors. One important holographic feature for these toy models is that, the entanglement entropy is given by the minimal number of cuts along a homologous surface in the network, which matches the area of the minimal surface of the celebrated Ryu-Takayanagi (RT) formula Ryu:2006bv ; Hubeny:2007xt ; Casini_2011 ; Lewkowycz:2013nqa ; Faulkner:2013ana ; Engelhardt:2014gca .
@@
Nevertheless, in most of these toy models their structures of the tensor network are usually discrete and only capture the features of the background geometry in a qualitative way. There is a big gap between number of cuts in a discrete network and the area of a surface in a smooth Riemannian manifold. Recently, a continuous network as a perfect tessellation of the background AdS geometry was constructed using the so-called partial-entanglement-entropy (PEE) threads Lin:2023rxc ; Lin:2024dho , which are bulk geodesics with a particular density distribution determined by the PEE structure of the boundary CFT. In such a network, the number of cuts along a surface exactly reproduces the area of this surface. In this work, we will develop two toy models of tensor network based on the PEE threads (or the PEE tensor network for short) in $d$-dimensional Poincaré AdS space. The first one is the factorized PEE tensor network constructed via tensor product of EPR pairs. It enjoys similar properties to the Harlow-Pastawski-Preskill-Yoshida (HaPPY) code Pastawski:2015qua which is built from perfect tensors, and perfectly reproduces the RT formula for spherical regions in the vacuum CFT state. In the second model, following Hayden:2016cfa ; Vasseur:2018gfy ; Jia:2020etj ; Cheng:2022ori ; Dong:2023kyr ; Kaya:2025vof we associate a random state to each bulk site to build the random PEE tensor network, where we can exactly reproduce the RT formula for generic boundary regions.
2 The network of PEE threads
@@
The partial entanglement entropy (PEE) Wen:2018whg ; Wen:2019iyq ; Wen:2020ech ; Han:2019scu ; Han:2021ycp ; Kudler-Flam:2019oru is a measure of entanglement $\mathcal{I}(A,B)$ between any two non-overlapping regions $A$ and $B$ featured by the key property of additivity. It can be determined by the following physical requirements Wen:2019iyq ; Vidal:2014aal :
@@
Here the regions are all non-overlapping, $A\cup\bar{A}$ makes a pure state and $S_{A}$ denotes the entanglement entropy of $A$. The solutions for these requirements were first studied in Casini_2011 as the mutual information that is additive in some special theories, and then the solutions are understood as the PEE in Wen:2019iyq , which is in general not a mutual information. Later it was gradually realized that, the requirement of normalization should not be imposed for an arbitrary region. In the case of the vacuum CFT state on a plane, the solution only exists if we impose the normalization requirement to spherical regions Wen:2019iyq . See also Vidal:2014aal ; Wen:2018whg ; Wen:2018mev ; Wen:2019iyq ; Kudler-Flam:2019oru for other prescriptions to construct the PEE which satisfy the above requirements, and see Basu:2023wmv ; Rolph:2021nan ; Lin:2022aqf ; Lin:2023orb ; Wen:2021qgx ; Wen:2022jxr ; Wen:2024uwr ; Camargo:2022mme for related discussions on PEE.
Due to the properties of additivity and permutation symmetry, any PEE $\mathcal{I}(A,B)$ can be further decomposed into a class of PEE between two different sites $\mathbf{x}$ and $\mathbf{y}$, or the two-point PEE $\mathcal{I}\left(\mathbf{x},\mathbf{y}\right)$, in the following way Wen:2019iyq ; Lin:2023rxc :
@@
Moreover, in the vacuum CFTd on a plane the two-point PEE can be fully determined by imposing all the above requirements for the spherical regions Lin:2023rxc ; Lin:2024dho , i.e.
@@
In the context of AdS/CFT, the vacuum CFTd state on a plane is dual to the Poincaré AdSd+1 space. The two-point PEE $\mathcal{I}\left(\mathbf{x},\mathbf{y}\right)$ was also proposed to have a natural holographic dual Lin:2023rxc ; Lin:2024dho , which is the density of the bulk geodesics connecting the boundary sites $\mathbf{x}$ and $\mathbf{y}$. The bulk geodesics whose density are determined in such a way are denoted as the PEE threads. Let us set that each PEE thread gives unit contribution to the PEE, then (1) captures the number of PEE threads connecting the two boundary regions $A$ and $B$. The collection of all the PEE threads forms a continuous, dense network permeating the entire time slice of the Poincaré AdSd+1, which we call the PEE network (see an example of $d=2$ in Fig.1). Remarkably, it was found that Lin:2023rxc ; Lin:2024dho the PEE network is a tessellation of the static AdS space in the following sense:
@@
One profound implication of the above prescription is that, minimizing the area of a surface $\Sigma_{A}$ homologous to any boundary region $A$ is equivalent to minimizing the number of intersections between $\Sigma_{A}$ (see the red line in Fig.1) and the PEE network. This further implies that we can exactly reproduce the Ryu-Takayanagi (RT) formula for holographic entanglement entropy by developing proper tensor network models where the entanglement entropy of a boundary region is captured by the minimal number of intersections between a homologous surface and the PEE network Lin:2023rxc ; Lin:2024dho ; Wen:2024uwr . Briefly speaking, the RT formula for the boundary region $A$ can be reformulated as follows,
@@
Figure 1: Extracted from Wen:2024uwr . The PEE network as a tessellation of a time slice of AdS3. Here the blue lines represent the PEE threads, the red line is a surface $\Sigma_{A}$ homologous to the boundary region AA.
3 Geometric setup for bulk tensors
In a model of tensor network, a vertex at site $\mathbf{z}$ with a set of legs corresponds to a quantum state represented by a local tensor $\mathcal{T}_{u_{1}u_{2}\cdots}(\mathbf{z})$, where the index $u_{i}$ characterizes the leg that is associated to a Hilbert space $\mathcal{H}_{i}$ spanned by the basis vectors $\{|u_{i}\rangle\}$. Then the quantum state in the Hilbert space $\bigotimes_{i}\mathcal{H}_{i}$ settled at $\mathbf{z}$ is defined by (assuming Einstein summation hereafter),
@@ 3 Geometric setup for bulk tensors
Figure 2: A tensor state located at the bulk site $\mathbf{z}$ (hollow white dot in the center). The legs connecting $\mathbf{z}$ is actually continuously distributed, which can be described by a vector field as in Lin:2023rxc . Only six legs are exhibited, while other legs are collectively represented by the blue region. The contraction between tensors at $\mathbf{z}$ and at adjacent site $\mathbf{z^{\prime}}$ is realized by projecting the total state into a maximally entangled state $|\mathbf{z}\mathbf{z^{\prime}}\rangle$ along the PEE thread in between.
@@
4 Factorized PEE tensor network
Let us consider a pair of PEE threads which makes a complete geodesic passing through $\mathbf{z}$, and we denote it as $C_{\mathbf{z}}(\mathbf{x},\mathbf{x^{\prime}})\equiv C_{\mathbf{z}}(\mathbf{x})\cup C_{\mathbf{z}}(\mathbf{x^{\prime}})$. Correspondingly we can divide the legs of $\mathbf{z}$ into pairs and rewrite the quantum state (8) as,
@@ 4 Factorized PEE tensor network
It has been demonstrated that Lin:2024dho , given a static boundary interval $A$ (or any spherical region in higher dimensions) and the corresponding RT surface $\gamma_{A}$, the PEE threads in $\mathcal{C}^{A\bar{A}}$ are the only threads that intersect with $\gamma_{A}$, and each thread only intersects once. This means the number of intersections between $\gamma_{A}$ and the PEE network is exactly $\mathcal{I}(A,\bar{A})$. Since the area of $\gamma_{A}$ can be computed by counting the number of intersections (3), we have
@@
which reproduces the RT formula for any static boundary single intervals and spherical regions. See the appendix A for more details for the bulk geometric picture, and another derivation for the RT formula by constructing isometric mappings from $\gamma_{A}$ to the boundary.
Nevertheless, for disconnected and non-spherical connected boundary regions, the homologous surfaces $\Sigma_{A}$ that only intersect with the threads in $\mathcal{C}^{A\bar{A}}$ no longer exists. Furthermore, the RT surface $\gamma_{A}$ intersects not only with all the threads in $\mathcal{C}^{A\bar{A}}$, but also with a subset of the threads in $\mathcal{C}^{A}$ or $\mathcal{C}^{\bar{A}}$ Lin:2024dho , which means $\mathcal{I}(A,\bar{A})<\frac{Area(\gamma_{A})}{4G}$. For these reasons we fail to reproduce the RT formula for generic boundary regions in this model.
5 Random PEE tensor network
In the second model, we also start from the PEE network, but set the states $|\mathcal{T}(\mathbf{z})\rangle$ (8) on the bulk vertices to be unit random states following Hayden:2016cfa . We will take the random average of such states when computing the entropy quantities. To be specific, we take an arbitrary reference state $|0_{\mathbf{z}}\rangle$ and define $|\mathcal{T}(\mathbf{z})\rangle=U_{\mathbf{z}}|0_{\mathbf{z}}\rangle$ where $U_{\mathbf{z}}$ is a unitary operator, then the random average is equivalent to an integration over $U_{\mathbf{z}}$ according to the Haar probability measure. Again, after contracting all the internal legs of adjacent bulk sites along all the PEE threads using (9), we obtain the random PEE tensor network representing a boundary state for the boundary open legs.
In Hayden:2016cfa , the authors have shown that in a general graph of network with random tensors on the vertices, if we take the large $v$ limit for the bond dimension of legs, the entanglement entropy for an arbitrary boundary region $A$ is dominant by the area term:
@@ 5 Random PEE tensor network
6 Discussions
In this work, we assign quantum states to the vertices in the PEE network of AdS space and build two PEE tensor network models, where the entanglement entropy can be explicitly computed. In our models, counting the minimal number of cuts in the tensor network is exactly computing the area of the RT surface in the semi-classical geometry of gravity. This provides the first example of taking natural continuous limit for a discrete tensor network to perfectly match the geometric background.
In the factorized model, the RT formula is only understood for spherical regions. This is not surprising as in this model all the PEE threads are decoupled from each other, which means the model is oversimplified for AdS/CFT. The factorized model reminds us of the bit thread configurations Freedman:2016zud ; Headrick:2017ucz , and we recommend the readers to consult Lin:2023rxc for their similarities and differences  [^2] In Lin:2023rxc it was demonstrated that bit threads configuration for spherical regions $A$ can be reproduced by the superposition of the vector fields describing the PEE threads emanating from $A$..
More interesting models can be developed by including coupling among the PEE threads. In the appendix B, besides regarding the tensors as tensor product of two-index tensors, we also construct a HaPPY-like Pastawski:2015qua PEE tensor network, where we associate a perfect tensor to each site in the PEE network. The RT formula for connected regions in AdS3 is derived using a similar greedy algorithm developed in Pastawski:2015qua for our HaPPY-like PEE tensor network.
As in Hayden:2016cfa ; Pastawski:2015qua , we can also set open legs for the bulk sites to built a holographic code mapping the bulk states to the boundary states, then we can study the bulk quantum correction to the holographic entanglement entropy and identify the analogue of the quantum extremal surface Engelhardt:2014gca . Perhaps the most interesting future direction is to generalize the PEE tensor network to geometric background beyond the Poincaré AdS space, thus construct a new framework to study holographic correspondence beyond AdS/CFT. This is promising because the Crofton formula works for more generic Riemann manifolds, and we can build the network of all the geodesics in the Kinematic Space. The generalization will be non-trivial, as it is quite possible that the geodesics in the bulk will not always anchor on the boundary, which may correspond to spaces containing black holes or singularities.
Acknowledgements.
The authors are supported by the NSFC Grant No. 12447108 and the Shing-Tung Yau Center of Southeast University. We thank Debarshi Basu for helpful discussions and related collaborations.